# Universal Beta Splatting

### v1

@@ 5.2 Comparison with the State of the Art
STATIC SCENES
NeRF Synthetic
chair 35.60 0.988 0.010 4.8 35.55 0.986 0.011 19.6 36.72 0.990 0.009 8.7
@@
DYNAMIC SCENES
7DGS-PBR
heart1 27.23 0.949 0.046 103.0 34.66 0.983 0.023 114.2 37.44 0.990 0.013 38.7
@@
Overall, UBS-6D demonstrates exceptional performance on challenging view-dependent scenarios in 6DGS-PBR. In addition, UBS-6D establishes new state-of-the-art results on NeRF Synthetic and Mip-NeRF 360 shown in Table 3 in Appendix.
@@
The consistent improvements across PSNR, SSIM, and LPIPS metrics confirm that UBS’s enhanced representational capacity translates to comprehensive quality gains. Figure 3 demonstrates that UBS achieves better visual quality in both static and dynamic scenes compared with baseline methods.
Training time. Beyond reconstruction quality, UBS demonstrates superior computational efficiency through its streamlined primitive design and optimized CUDA implementation. For static scenes, UBS-6D achieves significantly faster training than 6DGS on NeRF Synthetic (8.7 vs 12.7 minutes, 31.5% reduction) and Mip-NeRF 360 (25.2 vs 81.9 minutes, 69.2% reduction). On 6DGS-PBR, UBS-6D requires longer training time (27.7 vs 22.2 minutes) due to our use of higher primitive counts (300K vs 68K) to fully exploit Beta kernels’ representational capacity for complex volumetric effects. For dynamic scenes, UBS-7D delivers substantial training time reductions of 48.7% on 7DGS-PBR (59.9 vs 116.7 minutes) and 17.0% on D-NeRF (42.1 vs 50.7 minutes), both compared to 7DGS. These efficiency gains stem from our unified Beta kernel approach, which eliminates auxiliary spherical harmonic encodings while enabling efficient CUDA-accelerated conditional slicing operations.
@@ 5.3 Ablation Study
Spatial-Orthogonal Parameterization. Pure rotation-scale parameterization and full Cholesky without spatial orthogonality reduce quality by 0.29 dB and 0.56 dB, respectively, confirming the importance of preserving the spatial $\text{SO}(3)$ structure and maintaining flexibility of other dimensions. Beta-Modulated Conditioning. Removing Beta modulation decreases performance by 0.53 dB, with larger drops on dynamic scenes (lego: 1.25-1.25 dB), validating dimension-specific shape control for decoupling spatial-angular-temporal responses. Beta-Modulated Opacity. Disabling Beta-modulated opacity causes the most significant degradation (3.52-3.52 dB average), particularly on volumetric scenes (cloud: 6.68-6.68 dB), proving the product-form opacity crucial for anisotropic falloff. CUDA Acceleration. Custom CUDA kernels reduce training time by 58% (12.9 vs 30.7 minutes) and improve rendering speed by 26% (196.2 vs 155.8 FPS), enhancing real-time performance through fused Beta evaluation and conditional slicing operations.
@@ 6 Conclusion
Appendix A Comparison with State-of-the-art Methods
@@
Table 3: Comparison with SOTA methods on Mip-NeRF360 (Barron et al., 2022) and NeRF Synthetic (Mildenhall et al., 2020). We adopt the original number from corresponding papers when available. A dash (“-”) indicates results that were not reported.
@@
DBS (Liu et al., 2025) 28.60 0.844 0.182 34.64 0.973 0.028
@@
Appendix B Comparisons on the 6DGS-PBR dataset
Table 4 compares UBS-6D (Ours) with 3DGS (Kerbl et al., 2023), DBS (Liu et al., 2025), and 6DGS (Gao et al., 2025a) on the 6DGS-PBR dataset. UBS-6D consistently achieves the best performance across all seven scenes, surpassing all baselines in reconstruction accuracy and perceptual quality. On average, UBS-6D achieves 40.10 dB PSNR, 0.979 SSIM, and 0.039 LPIPS, outperforming 6DGS by +3.4 dB PSNR and DBS by +8.8 dB, with comparable training efficiency (27.7 min vs. 22.2/22.4 min). The clear gap between 6DGS/UBS-6D and 3DGS/DBS arises from the presence of conditional slicing, which allows the mean and covariance to adapt with viewing conditions. In contrast, 3DGS and DBS rely on static spatial kernels, limiting their ability to capture these view-dependent effects. By combining high-dimensional Beta kernels with conditional slicing, UBS-6D attains complex visual effects and photorealism while maintaining competitive runtime.
Table 4: Results on the 6DGS-PBR (Gao et al., 2025a) dataset. Train indicates time in minutes.
@@ Appendix B Comparisons on the 6DGS-PBR dataset
bunny 29.19 0.985 0.074 22.8 30.10 0.982 0.094 23.5 37.41 0.991 0.052 24.1 45.68 0.996 0.023 32.1
cloud 34.03 0.985 0.058 23.1 43.88 0.993 0.031 12.7 41.00 0.991 0.050 21.6 46.81 0.995 0.027 30.6
explosion 27.06 0.953 0.097 11.3 32.42 0.965 0.065 17.9 41.61 0.990 0.031 18.3 44.54 0.994 0.018 32.4
smoke 26.82 0.964 0.088 24.2 28.85 0.964 0.081 47.5 41.41 0.993 0.041 31.1 44.02 0.995 0.028 33.5
suzanne 22.45 0.885 0.159 19.5 26.58 0.923 0.113 30.3 26.96 0.928 0.106 32.9 27.35 0.931 0.098 35.1
dragon 26.53 0.812 0.123 12.0 29.71 0.844 0.093 20.7 34.95 0.933 0.044 16.3 38.06 0.974 0.034 23.7
ct-scan 25.69 0.917 0.099 4.5 27.52 0.923 0.091 4.3 33.46 0.964 0.058 11.0 34.24 0.969 0.046 6.5
avg 27.40 0.929 0.100 16.8 31.29 0.942 0.081 22.41 36.68 0.970 0.055 22.2 40.10 0.979 0.039 27.7
Appendix C Per-Primitive Parameter Analysis
@@
Appendix D Rotation Parameterization Analysis
To validate our spatial-orthogonal Cholesky parameterization design, we conduct a comprehensive comparison of different spatial rotation parameterizations for covariance matrix construction on the bicycle scene from the Mip-NeRF 360 dataset (Barron et al., 2022) using UBS-3D (pure spatial Beta splatting without view or temporal dimensions). Note that this analysis focuses solely on spatial rotation parameterization within the 3D spatial subspace, not the full $N$-dimensional covariance structure used in UBS-6D or UBS-7D.

### v2

@@ 5.2 Comparison with the State of the Art
STATIC SCENES NeRF Synthetic chair 35.60 0.988 0.010 4.8 35.55 0.986 0.011 19.6 36.72 0.990 0.009 8.7
@@
DYNAMIC SCENES 7DGS-PBR heart1 27.23 0.949 0.046 103.0 34.66 0.983 0.023 114.2 37.44 0.990 0.013 38.7
@@
Overall, UBS-6D demonstrates exceptional performance on challenging view-dependent scenarios in 6DGS-PBR. In addition, UBS-6D establishes new state-of-the-art results on NeRF Synthetic and Mip-NeRF 360 shown in Table 4 in Appendix A.
@@
The consistent improvements across PSNR, SSIM, and LPIPS metrics confirm that UBS’s enhanced representational capacity translates to comprehensive quality gains. Figure 3 demonstrates that UBS achieves better visual quality in both static and dynamic scenes compared with baseline methods. Additional quantitative benchmark comparisons are provided in Appendix A.
Efficiency Analysis. Beyond reconstruction quality, UBS demonstrates superior computational efficiency through its streamlined primitive design and optimized CUDA implementation. Table 2 summarizes the efficiency comparison across all benchmarks, with detailed per-scene results provided in Table 8 in the Appendix B. For static scenes, UBS-6D achieves significantly faster training than 6DGS on NeRF Synthetic (8.7 vs 12.7 minutes, 31.5% reduction) and Mip-NeRF 360 (25.2 vs 81.9 minutes, 69.2% reduction) with competitive memory usage. On 6DGS-PBR, UBS-6D requires longer training time (27.7 vs 22.2 minutes) due to our use of higher primitive counts (300K vs 68K) to fully exploit Beta kernels’ representational capacity for complex volumetric effects. For dynamic scenes, UBS-7D delivers substantial training time reductions of 48.7% on 7DGS-PBR (59.9 vs 116.7 minutes) and 17.0% on D-NeRF (42.1 vs 50.7 minutes), both compared to 7DGS. These efficiency gains stem from our unified Beta kernel approach, which eliminates auxiliary spherical harmonic encodings while enabling efficient CUDA-accelerated conditional slicing operations.
Table 2: Efficiency comparison on static and dynamic benchmarks (averaged across scenes). ‘Train’ indicates training time in minutes; ‘Mem’ indicates storage size in MB. Detailed per-scene results and per-primitive parameter comparisons are provided in Appendix B.
Dataset 3DGS/4DGS 6DGS/7DGS UBS (Ours)
[…]
NeRF Synthetic 695.6 289.3K 68.5 3.6 648.0 240.7K 69.8 12.7 345.7 300.0K 40.1 8.7
Mip-NeRF 360 178.7 3.3M 775.9 21.7 157.4 2.0M 581.2 81.9 100.4 3.1M 409.0 25.2
6DGS-PBR 305.9 147.7K 34.9 16.8 339.8 84.3K 24.4 22.2 290.2 300.0K 40.1 27.7
7DGS-PBR 192.6 642.0K 613.1 127.2 376.0 88.4K 29.7 116.7 359.6 234.3K 39.4 59.9
D-NeRF 296.4 255.3K 380.3 52.6 377.8 43.8K 16.0 50.7 368.2 168.8K 28.4 42.1
@@ 5.3 Ablation Study
Spatial-Orthogonal Parameterization. Pure rotation-scale parameterization and full Cholesky without spatial orthogonality reduce quality by 0.29 dB and 0.56 dB, respectively, confirming the importance of preserving the spatial $\text{SO}(3)$ structure and maintaining flexibility of other dimensions. Beta-Modulated Conditioning. Removing Beta modulation decreases performance by 0.53 dB, with larger drops on dynamic scenes (lego: 1.25-1.25 dB), validating dimension-specific shape control for decoupling spatial-angular-temporal responses. Beta-Modulated Opacity. Disabling Beta-modulated opacity causes the most significant degradation (3.52-3.52 dB average), particularly on volumetric scenes (cloud: 6.68-6.68 dB), proving the gated product form opacity crucial for anisotropic falloff. CUDA Acceleration. Custom CUDA kernels reduce training time by 58% (12.9 vs 30.7 minutes) and improve rendering speed by 26% (196.2 vs 155.8 FPS), enhancing real-time performance through fused Beta evaluation and conditional slicing operations.
@@ 6 Conclusion
Appendix A More Comparisons
A.1 Comparison with State-of-the-art Methods
@@
Table 4: Comparisonh with SOTA methods on Mip-NeRF360 (Barron et al., 2022) and NeRF Synthetic (Mildenhall et al., 2020).
@@
DRKS (Huang et al., 2025) 26.76 0.787 0.236 33.82 - -
Textured GS (Chao et al., 2025) 27.35 0.827 0.186 33.24 0.967 0.043
Quadratic GS (Zhang et al., 2025) 27.39 0.813 0.213 - - -
Disc-GS (Qu et al., 2024) 28.01 0.833 0.189 - - -
DBS (Liu et al., 2025a) 28.60 0.844 0.182 34.64 0.973 0.028
@@
Table 5 reports a quantitative comparison of UBS-7D against state-of-the-art kernel-based methods on the real-world N3V dataset (Li et al., 2022). UBS-7D achieves the best overall average PSNR of 32.22 dB, outperforming Ex4DGS (32.11 dB) and 7DGS (31.59 dB), while winning in 4 out of 6 scenes. This demonstrates the unified Beta primitive’s effectiveness in capturing complex temporal dynamics and appearance variations in real-world dynamic content. Earlier methods 4DGS (30.19 dB) and 4DGaussians (28.63 dB) show significantly larger performance gaps, validating the advantage of our N-D anisotropic kernel design for challenging 4D reconstruction tasks.
Table 5: Comparison with other kernel methods on the real-world N3V dataset (Li et al., 2022).
Model Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak Average
4DGS (Yang et al., 2024a) 26.51 32.11 31.74 26.93 31.44 32.42 30.19
4DGaussians (Wu et al., 2024) 26.69 31.89 25.88 27.54 28.07 31.73 28.63
Ex4DGS (Lee et al., 2024b) 28.79 33.23 33.73 29.29 33.91 33.69 32.11
7DGS (Gao et al., 2025b) 29.07 32.23 32.93 28.83 33.07 33.41 31.59
UBS-7D (Ours) 29.73 33.30 33.88 28.77 33.83 33.83 32.22
A.2 Comparisons on the 6DGS-PBR dataset
Table 6 compares UBS-6D (Ours) with 3DGS (Kerbl et al., 2023), DBS (Liu et al., 2025a), and 6DGS (Gao et al., 2025a) on the 6DGS-PBR dataset. UBS-6D consistently achieves the best performance across all seven scenes, surpassing all baselines in reconstruction accuracy and perceptual quality. On average, UBS-6D achieves 40.10 dB PSNR, 0.979 SSIM, and 0.039 LPIPS, outperforming 6DGS by +3.4 dB PSNR and DBS by +8.8 dB, with comparable training efficiency (27.7 min vs. 22.2/22.4 min). The clear gap between 6DGS/UBS-6D and 3DGS/DBS arises from the presence of conditional slicing, which allows the mean and covariance to adapt with viewing conditions. In contrast, 3DGS and DBS rely on static spatial kernels, limiting their ability to capture these view-dependent effects. By combining high-dimensional Beta kernels with conditional slicing, UBS-6D attains complex visual effects and photorealism while maintaining competitive runtime.
Table 6: Results on the 6DGS-PBR (Gao et al., 2025a) dataset. ‘Train’ indicates time in minutes.
@@ A.2 Comparisons on the 6DGS-PBR dataset
bunny 29.19 0.985 22.8 30.10 0.982 23.5 37.41 0.991 24.1 45.68 0.996 32.1
cloud 34.03 0.985 23.1 43.88 0.993 12.7 41.00 0.991 21.6 46.81 0.995 30.6
explosion 27.06 0.953 11.3 32.42 0.965 17.9 41.61 0.990 18.3 44.54 0.994 32.4
smoke 26.82 0.964 24.2 28.85 0.964 47.5 41.41 0.993 31.1 44.02 0.995 33.5
suzanne 22.45 0.885 19.5 26.58 0.923 30.3 26.96 0.928 32.9 27.35 0.931 35.1
dragon 26.53 0.812 12.0 29.71 0.844 20.7 34.95 0.933 16.3 38.06 0.974 23.7
ct-scan 25.69 0.917 4.5 27.52 0.923 4.3 33.46 0.964 11.0 34.24 0.969 6.5
avg 27.40 0.929 16.8 31.29 0.942 22.41 36.68 0.970 22.2 40.10 0.979 27.7
Appendix B Efficiency Analysis
B.1 Per-Primitive Parameter Analysis
@@
B.2 Computational Efficiency
Table 8: Efficiency comparison with 3DGS and 6DGS on static benchmarks of NeRF Synthetic (Mildenhall et al., 2020), Mip-NeRF (Barron et al., 2022), 6DGS-PBR (Gao et al., 2025a), and with 4DGS and 7DGS on dynamic benchmarks of 7DGS-PBR (Gao et al., 2025b) and D-NeRF (Pumarola et al., 2021). ‘Train’ means training time in minutes; ‘Mem’ means storage size in MB.
Dataset Scene 3DGS 6DGS UBS-6D (Ours)
[…]
STATIC SCENES NeRF Synthetic chair 412.8 489.0K 115.7 4.8 420.4 410.7K 119.1 19.6 356.3 300.0K 40.1 8.7
drums 562.3 390.9K 92.5 4.1 555.8 317.2K 92.0 15.4 349.1 300.0K 40.1 8.1
ficus 728.5 264.9K 62.7 3.0 757.4 226.4K 65.6 11.3 356.6 300.0K 40.1 8.2
hotdog 831.5 187.8K 44.4 3.5 700.5 161.0K 46.7 10.0 352.5 300.0K 40.1 9.1
lego 582.4 347.0K 82.1 4.0 247.7 273.3K 79.2 15.3 340.3 300.0K 40.1 8.4
materials 1286.1 160.1K 37.9 2.7 1098.2 125.0K 36.2 7.8 360.7 300.0K 40.1 8.9
mic 614.9 195.7K 46.3 3.0 756.9 174.3K 50.5 9.4 326.1 300.0K 40.1 9.5
ship 546.4 279.2K 66.0 4.1 646.8 237.9K 69.0 12.7 330.3 300.0K 40.1 8.8
avg 695.6 289.3K 68.5 3.6 648.0 240.7K 69.8 12.7 345.7 300.0K 40.1 8.7
Mip-NeRF 360 bicycle 105.2 6.0M 1421.5 29.2 105.2 3.3M 957.2 114.9 43.9 6.0M 801.1 30.2
flowers 150.9 3.6M 850.0 19.6 150.9 2.4M 707.2 88.8 84.3 3.0M 400.5 19.6
garden 84.2 5.7M 1358.3 30.4 84.2 4.2M 1219.4 164.9 52.9 5.0M 667.6 36.8
stump 151.2 4.6M 1095.1 22.7 190.5 1.7M 488.2 60.2 64.9 4.0M 536.3 18.9
treehill 161.6 3.8M 886.5 20.1 130.9 2.6M 755.8 94.0 72.4 3.5M 467.3 21.9
room 222.0 1.6M 369.1 18.2 171.2 1.1M 314.8 51.4 147.2 1.5M 200.3 21.9
counter 233.5 1.2M 281.4 17.6 203.7 0.8M 239.8 43.5 143.8 1.5M 200.3 27.8
kitchen 183.0 1.8M 424.7 21.5 147.4 1.0M 285.9 71.3 144.8 1.5M 200.3 26.6
bonsai 317.1 1.3M 296.6 15.7 233.4 0.9M 262.7 48.4 147.1 1.5M 200.3 22.7
avg 178.7 3.3M 775.9 21.7 157.4 2.0M 581.2 81.9 100.4 3.1M 409.0 25.2
6DGS-PBR bunny 123.8 79.8K 18.87 22.8 349.5 9.7K 2.8 24.1 257.9 300.0K 40.1 32.1
cloud 370.5 60.0K 14.19 23.1 429.8 17.9K 5.2 21.6 247.7 300.0K 40.1 30.6
explosion 400.3 51.2K 12.11 11.3 302.8 25.9K 7.5 18.3 275.6 300.0K 40.1 32.4
smoke 176.6 55.8K 13.20 24.2 207.7 21.0K 6.1 31.1 225.0 300.0K 40.1 33.5
suzanne 186.5 316.2K 74.78 19.5 94.9 208.3K 60.4 32.9 325.9 300.0K 40.1 35.1
dragon 275.9 241.4K 57.09 12.0 266.9 127.8K 37.1 16.3 334.9 300.0K 40.1 23.7
ct-scan 607.2 229.2K 54.20 4.5 726.8 179.3K 52.0 11.0 370.2 300.0K 40.1 6.5
avg 305.9 147.7K 34.9 16.8 339.8 84.3K 24.4 22.2 290.2 300.0K 40.1 27.7
Dataset Scene 4DGS 7DGS UBS-7D (Ours)
[…]
DYNAMIC SCENES 7DGS-PBR heart1 186.9 694.0K 444.6 103.0 401.0 82.4K 27.7 114.2 244.6 328.0K 55.1 38.7
heart2 160.4 869.2K 556.9 103.4 384.6 101.5K 34.1 384.6 297.1 328.0K 55.1 33.5
cloud 219.0 216.9K 138.9 123.7 386.2 44.2K 14.8 102.6 473.6 150.0K 25.2 50.0
dust 296.1 357.7K 229.2 97.0 394.8 10.9K 3.7 69.8 406.0 150.0K 25.2 46.5
flame 151.2 947.8K 607.2 113.7 371.6 15.1K 5.1 74.1 461.4 150.0K 25.2 53.7
suzanne 141.8 766.1K 1,701.6 222.5 317.9 276.3K 92.8 193.9 275.0 300.0K 50.4 136.8
avg 192.6 642.0K 613.1 127.2 376.0 88.4K 29.7 116.7 359.6 234.3K 39.4 59.9
D-NeRF b.balls 219.9 276.1K 277.7 50.7 213.1 127.4K 43.5 86.6 366.6 150.0K 25.2 49.6
h.warrior 299.4 298.4K 212.6 35.3 431.5 8.7K 2.9 31.3 395.2 150.0K 25.2 36.3
hook 325.1 174.7K 183.3 38.0 432.8 21.7K 7.3 35.7 452.2 150.0K 25.2 34.0
j.jacks 366.0 143.7K 534.8 66.9 432.8 21.7K 5.3 34.1 376.4 150.0K 25.2 43.2
lego 320.0 186.2K 355.0 55.4 365.0 68.6K 31.4 78.5 347.4 150.0K 25.2 31.3
mutant 341.6 138.7K 88.9 39.1 395.8 33.9K 11.5 42.5 258.7 300.0K 50.4 60.0
standup 330.4 142.5K 107.1 34.4 399.2 12.7K 4.3 33.5 389.4 150.0K 25.2 41.1
trex 169.2 682.4K 1,085.2 100.7 352.4 62.0K 21.4 63.4 359.4 150.0K 25.2 41.6
avg 296.4 255.3K 380.3 52.6 377.8 43.8K 16.0 50.7 368.2 168.8K 28.4 42.1
Table 9: Impact of primitive capacity on NeRF Synthetic dataset. ‘Mem’ means storage size in MB.
Scene 100K Primitives 200K Primitives 300K Primitives
[…]
chair 36.16 0.989 13.4 543.4 36.62 0.990 26.7 341.3 36.62 0.990 40.1 356.3
drums 26.88 0.959 13.4 529.3 27.01 0.959 26.7 376.3 27.03 0.959 40.1 349.1
ficus 36.17 0.988 13.4 574.4 36.44 0.989 26.7 431.9 36.57 0.989 40.1 356.6
hotdog 38.68 0.988 13.4 534.7 38.81 0.988 26.7 366.9 38.94 0.988 40.1 352.5
lego 35.66 0.981 13.4 572.1 36.67 0.984 26.7 370.7 36.97 0.985 40.1 340.3
materials 32.49 0.974 13.4 531.7 32.85 0.976 26.7 367.7 32.93 0.977 40.1 360.7
mic 37.48 0.994 13.4 493.1 37.79 0.994 26.7 346.7 37.92 0.994 40.1 326.1
ship 31.88 0.911 13.4 543.4 32.10 0.913 26.7 344.8 32.09 0.913 40.1 330.3
avg 34.42 0.973 13.4 540.3 34.79 0.974 26.7 368.3 34.89 0.974 40.1 345.7
Table 8 presents a comprehensive efficiency comparison between UBS and baseline methods. For static scenes, UBS-6D achieves 31% faster training than 6DGS on NeRF Synthetic (8.7 vs 12.7 minutes) with 43% memory reduction (40.1 vs 69.8 MB) and 345.7 FPS rendering. On Mip-NeRF 360, UBS achieves the lowest memory usage (409.0 MB, 47% and 30% reduction vs 3DGS and 6DGS) with 100.4 FPS. On 6DGS-PBR, UBS achieves comparable training times (27.7 vs 22.2 minutes for 6DGS) with 290.2 FPS. For dynamic scenes, UBS-7D reduces training time by 53% and 49% compared to 4DGS and 7DGS on 7DGS-PBR (59.9 vs 127.2 and 116.7 minutes) with 359.6 FPS, and by 20% and 17% on D-NeRF (42.1 vs 52.6 and 50.7 minutes) with 368.2 FPS. The efficiency gains stem from UBS’s unified Beta kernel design, which eliminates auxiliary spherical harmonic encodings. Moreover, unlike previous Gaussian methods that rely on heuristic densification and pruning, UBS employs MCMC-based optimization that allows direct control over primitive count, enabling flexible speed-quality trade-offs as demonstrated in Table 9.
Table 9 demonstrates this flexibility by analyzing the speed-quality trade-off across varying primitive capacities (100K, 200K, 300K) on NeRF Synthetic. Results reveal that reducing primitives from 300K to 100K decreases quality by only 0.47 dB PSNR (34.89 to 34.42 dB) while significantly improving rendering speed by 56% (345.7 to 540.3 FPS) and reducing memory by 67% (40.1 to 13.4 MB). This demonstrates that UBS can be optimized for speed-critical applications by adaptively adjusting primitive counts to balance quality and efficiency.
Appendix C Spatial-Orthogonal Cholesky Parameterization